A system for CS4 administration

The current situation

Towards the end of each academic year, the Syllabus Committee determines which courses will be available to CS4 students in the following year.

At the end of each academic year, the Head of Department allocates duties to members of teaching staff and others; in particular, one person is assigned to lecture each of the courses which are supposed to be available in the following year. (We’ll call these people lecturers for simplicity.)

Each lecturer updates the CS4 Course Handbook entry for his or her course. The CS4 co-ordinator updates other parts of the handbook, and checks the course entries produced by the lecturers.

Somebody in the Undergraduate Teaching Office (from now on we’ll call any such person “the UTO”) produces the paper version of the course handbook; the CS4 co-ordinator produces the HTML version by running latex2html on the latex source.

The CS3 Co-ordinator is supposed to give a list of the students entering CS4 from CS3 both to the CS4 co-ordinator, and to the UTO. The CS4 co-ordinator tells the UTO about any students entering CS4 other than from CS3, for example, non-graduating students. The UTO keeps the master list of all CS4 students, and updates the mailing list cs4.

CS4 students provisionally register for courses by filling in paper forms and handing them in to the Undergraduate Teaching Office. The UTO checks that every student who registers is listed as a CS4 student, and that every CS4 student is registered for a reasonable number for courses.

The UTO then produces for lecturers lists of the students taking their courses. These arrive not sooner than week 3, in particular, too late to be useful for letting lecturers know how many copies of things to make...

The investigation

The Department has asked you to investigate the possibility of developing a system to automate parts of this process, because they hope it may be possible to:

• decrease the burden of routine work on all staff, especially the poor overworked CS4 co-ordinator

• allow students to register for courses on-line

• make it easy to obtain (from the UTO) up-to-date, reliable information;

• improve the traceability of such information;

• make information such as the course handbook and lists of students taking courses available sooner.

Draw a simple use case diagram for such a system.

Hint: not all the information you’ve been given will be relevant at this point.